OpenStudy (anonymous):

A bouncing ball reaches a height of 54 inches at its first peak, 36 inches at its second peak, and 24 inches at its third peak. Which formula represents this scenario?

OpenStudy (anonymous):

bx

OpenStudy (anonymous):

A) f(x) = 54[2/3]x B) f(x) = 54[2/3]x-1 C) f(x) =2/3(54)x D) f(x) = 2/3(54)x-1

Mercury:

Asked & answered here: https://questioncove.com/study#/updates/5ef15664405cc5db8ab66800